OpenStudy (anonymous):

How do I write the series in summation notation for 1-3+9-27+81......

OpenStudy (anonymous):

Do you know the pattern?

OpenStudy (anonymous):

hint: \[\sum_{i=1}^{n} a^i......\]

OpenStudy (anonymous):

\[a _{i}\] actually. To find \[a _{i}\] identify a pattern.

OpenStudy (anonymous):

no its a^i

OpenStudy (anonymous):

the index here is an exponent

OpenStudy (anonymous):

It's geometric I believe ?

OpenStudy (anonymous):

yes and it takes advantage of the fact that a negative times a negative equals a positive

OpenStudy (anonymous):

Yes , so now I'm stuck on how to solve it ... It's confusing at least right now it is

OpenStudy (anonymous):

well its easy to see that we are involved with powers of 3 here...correct?

OpenStudy (anonymous):

\[\sum_{i = 1}^{\infty}\] This is what we know so far

OpenStudy (anonymous):

Yes ....

OpenStudy (anonymous):

We are trying to find the formula, to do this we must identify a pattern

OpenStudy (anonymous):

\[\sum_{i=1}^{n} -3^i\]

OpenStudy (anonymous):

check that and make sure you completely understand why it is correct

OpenStudy (anonymous):

Pattern as in r= -3 right /; ?

OpenStudy (anonymous):

Oh so I am correct ?

OpenStudy (anonymous):

Yes, -3 but you multiply.

OpenStudy (anonymous):

To what exactly ?

OpenStudy (anonymous):

you need to read more about summation (sigma) notation. heres a good site http://tutorial.math.lamar.edu/Classes/CalcI/SummationNotation.aspx

OpenStudy (anonymous):

If it was \[-3^{i}\] then then \[a_{2}\] would be -9 and this is not correct.

OpenStudy (anonymous):

no because -3*-3 =9

OpenStudy (anonymous):

\[\sum_{i=1}^{n} (-3)^i\]

OpenStudy (anonymous):

Yes it works with ()

OpenStudy (anonymous):

Sorry you are right it is exponential.

OpenStudy (anonymous):

Ok thanks & for the site too , @Edutopia & @af0998

OpenStudy (anonymous):

np here is another site that explains things simpler http://www.themathpage.com/aprecalc/sigma.htm

OpenStudy (anonymous):

Sum is S=1-3+9-27+..... or S=(1+9+81+....)-(3+27+243+....) or S=(a geometric series with a=1,r=9 upto n terms)-(a geometric series with a=3,r=9 upto n terms). or \[S=\sum_{n=0}^{n} 3^{2n}-\sum_{n=0}^{n}3^{2n+1}\] Another method: as suggested by @Edutopia